OpenStudy (anonymous):

Rewrite this in slope intercept form, x – y = -3

OpenStudy (saifoo.khan):

the standard form is: y = mx + b y = x + 3

OpenStudy (anonymous):

ty sensei -bows lol

OpenStudy (saifoo.khan):

\[\Huge \begin{array}l\color{red}{\text{L}}\color{orange}{\text{O}}\color{#9c9a2e}{\text{L}}\color{green}{\text{,}}\color{blue}{\text{ }}\color{purple}{\text{N}}\color{purple}{\text{o}}\color{red}{\text{ }}\color{orange}{\text{P}}\color{#9c9a2e}{\text{r}}\color{green}{\text{o}}\color{blue}{\text{b}}\color{purple}{\text{l}}\color{purple}{\text{e}}\color{red}{\text{m}}\color{orange}{\text{!}}\color{#9c9a2e}{\text{ }}\color{green}{\text{=}}\color{blue}{\text{)}}\color{purple}{\text{}}\end{array}\]

OpenStudy (anonymous):

y = mx + b y = x + 3

jimthompson5910 (jim_thompson5910):

Two ways to do it: 1) x – y = -3 x = -3+y x + 3 = y y = x+3 OR 2) x – y = -3 – y = -3-x -y = -x-3 y = x+3 ... Multiply both sides by -1 to isolate y here In either case, we end up with y = x+3 which is in y = mx+b form where m=1 and b=3. So the slope is m = 1 and the y-intercept is (0,3)